OpenStudy (anonymous):

@DebbieG pleaseeeeee @Blaze @Osako heeelpp !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! 19^2+24^2=c^2 361+576=c^2 sqrt937=c^2sqrt

OpenStudy (anonymous):

What are you trying to find first?

OpenStudy (yamyam70):

How to solve a math problem :) # 1. Don't panic.

OpenStudy (anonymous):

What are you studying?

OpenStudy (anonymous):

im dras im drawing it

OpenStudy (anonymous):

If you are trying to find c by itself then: A. c=19+24= 43 B. c=sqrt(937) C. c=sqrt(937), since the sqrt cancels out the square (x^2)

OpenStudy (debbieg):

What is your question? What you have above is pretty much it, except that you need to remember that when you take the square root, you get two solutions: \(\Large x^2=a\) \(\Large \sqrt{x^2}=\pm \sqrt{a}\) \(\Large x= \sqrt{a}\) or \(\Large x= -\sqrt{a}\)

OpenStudy (debbieg):

I don't know what you mean above with your A, B, C @osako. A is certainly wrong. B & C are the same thing (?) and are basically correct, except again, you did not find both solutions, only the positive one.

OpenStudy (anonymous):

A B C were just to indicate the order so it doesn't look as if they were just one answer. I could of used 1 2 and 3 but since this is math I wanted to distinguish it from numbers.

OpenStudy (debbieg):

Uhm, ok, but again - they are wrong. Where in the world do you get (a) from?? c=19+24= 43 DOES NOT FOLLOW from 19^2+24^2=c^2 And b and c certainly don't follow from your (a) above; and are missing the \(\pm\). Just want to make sure that the person who asked the original problem isn't confused by that post.